The demand for good (Qx) is given by the following equation: Qx = 20,200 - 12.5 Px + 5 Py-M + 1.5 Ax Suppose the firm spends please calculate carefully
Answer #1

Ox=20200-12.58 tspy-mt Is Ax Ax= 3000 lx=80, Py=60 imo 22000 a) Elasticity with respect to Px = dax px dPX QX Qx = 20200 - 21PAG b) Demand is witaey elastic ») elocity तार कर Ox = 20,200 -12.5px +5 (60) 22000 +1:5(3000) 20,200 -12.5x + 300-22000 +45oPAGE Total Revenue (TR) = Px Qx = 2400 x-Qx² 12.5 MR = DIR Cox 240 2x 12.5 profit maximization of a monopolust - MRMC 60 =) 2